OpenStudy (anonymous):

4x+4y=52 and -4x+2y=-28 the answer is x=12 and y=4????

OpenStudy (anonymous):

4x+4(4)=52 4x+16=52 4x=52-16 4x=36 x=9

OpenStudy (anonymous):

adding both the equations, 4x+4y= 52 -4x+2y=-28 ------------ 0+6y=24 =>y=24/6 =>y=3 put y=3 in any of the eqn. for instance while putting in first eqn you get... 4x+4(3)=52 =>4x=52-12 =>4x=40 =>x=40/4 =>x=10 Inorder to verify your ans put both x and y value in any one of the equation and make sure that it satisfies as follows. 4(10)+4(3)=52 40+12=52 52=52 you can see both left hand side and right hand side are equal. hence satisfies so the ans is x=10,y=3

OpenStudy (anonymous):

y is actually 5

OpenStudy (anonymous):

If the value of y is given then go for substituting it in the eqn,but you will get 2 different ans as shown, 4x+4(5)=52 =>4x=32 =>x=8 also, -4x+2(5)=-28 =>-4x=-28+10 -4x=-18 =>x=18/4 If you're looking for a common soln for both eqn then it is x=10,y=3 Otherwise if you want solve for x for both the equation given the value of y as 5, you 2 different values of x